# Multi-Source Position and Direction-of-Arrival Estimation Based on Euclidean Distance Matrices

### v1

@@ Abstract
A popular method to estimate the positions or directions-of-arrival (DOAs) of multiple sound sources using an array of microphones is based on steered-response power (SRP) beamforming. For a three-dimensional scenario, SRP-based methods need to jointly optimize three continuous variables for position estimation or two continuous variables for DOA estimation. This can be computationally expensive, especially when high localization accuracy is desired. In this paper, we propose novel methods for multi-source position and DOA estimation by exploiting properties of Euclidean distance matrices (EDMs) and their respective Gram matrices. All methods require estimated time-differences of arrival (TDOAs) between the microphones. In the proposed multi-source position estimation method only a single continuous variable, representing the distance between each source and a reference microphone, needs to be optimized. For each source, the optimal continuous distance variable and set of candidate TDOA estimates are determined by minimizing a cost function that is defined using the eigenvalues of the Gram matrix. The estimated relative source positions are then mapped to estimated absolute source positions by solving an orthogonal Procrustes problem for each source. The proposed multi-source DOA estimation method entirely eliminates the need for continuous variable optimization by defining a relative coordinate system per source such that one of its coordinate axes is aligned with the respective source DOA. The optimal set of candidate TDOA estimates is determined by minimizing a cost function that is defined using the eigenvalues of a rank-reduced Gram matrix. The computational cost of the proposed EDM-based methods is significantly reduced compared to the SRP-based methods, and for EDM-based DOA estimation the need for continuous variable optimization is even entirely eliminated. For two sources in a noisy and reverberant environment, experimental results for different source and microphone configurations show that the proposed EDM-based method consistently outperforms the SRP-based method in terms of position and DOA estimation accuracy.
@@ I Introduction
Estimating the locations, i.e., positions or directions-of-arrival (DOAs), of multiple speech sources in a noisy and reverberant environment is important for several applications, such as source tracking, speech enhancement, and speaker extraction [1, 2, 3, 4]. In this paper, we will consider both compact microphone arrays with relatively small inter-microphone distances as well as spatially distributed microphones with large inter-microphone distances. For compact microphone arrays the sources will be assumed to be in the far field, such that we will focus on estimating the DOA (azimuth and elevation) of each source, whereas for spatially distributed microphones we will focus on estimating the three-dimensional position of each source.
@@
Existing model-based methods for source localization fall into two main categories [1, 2]: one-step methods such as steered-response power methods [5, 6, 7] or the subspace-based multiple signal classification (MUSIC) method [8], and two-step methods, which rely on the prior estimation of variables such as time-differences of arrival (TDOAs) between microphones [9]. In addition, recently several learning-based methods [10, 11, 12, 13, 14, 15, 16] for source localization have been proposed, showing promising results. However, most learning-based methods are trained for a specific array geometry and do not support source localization with arbitrary array geometries. In this paper, we only consider model-based localization methods, which provide the flexibility to use different array geometries without the need to retrain a neural network. A popular model-based method for single-source DOA or position estimation is based on the steered-response power with phase transform (SRP-PHAT) functional [6, 17, 18, 19, 20, 21, 22, 23]. While DOA estimation requires jointly optimizing the SRP-PHAT functional in up to two continuous variables (azimuth and elevation), position estimation requires jointly optimizing the SRP-PHAT functional in up to three continuous variables (x, y, and z coordinates). To perform accurate source localization, the SRP-PHAT functional needs to be evaluated on a discrete multi-dimensional grid with a high enough resolution in each variable resulting in a high computational complexity [21, 24]. To avoid the multi-dimensional search of SRP-based methods, in [25] we proposed a method to estimate the three-dimensional position of a single source based on Euclidean distance matrices. This method minimizes a cost function defined using the eigenvalues of a Gram matrix associated with an EDM containing the inter-microphone distances and the source-microphone distances. Using estimated TDOAs, this problem can be reformulated as an optimization in only a single distance variable, where the solution represents the distance between the source and the reference microphone. Several model-based approaches have also been proposed for multi-source DOA and position estimation, e.g., based on determining multiple peaks in the SRP-PHAT functional or MUSIC spectrum [8, 7, 26, 27, 28, 29]. It should however be noted that most of these approaches focus on multi-source DOA estimation and not on multi-source position estimation. Moreover, many approaches rely on data association/clustering of acoustic features [26, 30, 31] and are quite sensitive to reverberation.
@@
In this paper, we first extend the EDM-based single-source position estimation method in [25] to multi-source position estimation, assuming the number of sources $S$ to be known. For each combination of candidate TDOA estimates, the optimal distance variable is determined by minimizing the EDM-based cost function using a one-dimensional exhaustive search. Assuming that each source position corresponds to a unique combination of TDOAs, the optimal set of candidate TDOA estimates and corresponding source distances are determined, thereby also solving the association of candidate TDOA estimates to sources. The estimated relative position of each source to the microphone array can then be determined from the Gram matrix with the estimated source distance. The estimated relative positions are then mapped to estimated absolute source positions by solving an orthogonal Procrustes problem for each source. As the second contribution of this paper, we propose an EDM-based method to estimate the DOAs of multiple sources relative to a compact microphone array, which does not require continuous variable optimization. We construct a rank-reduced Gram matrix by subtracting a rank-1 matrix based on the estimated TDOAs from the Gram matrix associated with the EDM containing the microphone distances. We define a cost function using the eigenvalues of this rank-reduced Gram matrix and, similarly as for multi-source position estimation, determine the optimal set of candidate TDOA estimates by considering the $S$ smallest values of this cost function. The DOA of each source is then estimated from the mapping between the absolute coordinate system and the relative coordinate system, which depends on the combinations of TDOA estimates corresponding to each source.
@@
Two sets of experiments are conducted for several simulated scenarios with two sources and six microphones in a room with mild background noise and reverberation. The results of the first experiment considering spatially distributed microphones show that the proposed EDM-based method consistently outperforms the SRP-based method in estimating the of the sources for all considered positions. Similarly, the results of the second experiment considering compact microphone arrays show that the proposed EDM-based method also outperforms the SRP-based method in estimating the DOAs of the sources for all considered distances between the sources and the microphone array. Since the EDM-based methods reduce or eliminate the number of continuous variables that need to be optimized, the runtime can be considerably reduced compared to the SRP-based methods (about 380 times for position estimation and about 25 times for DOA estimation).
@@ II Theoretical Background
We consider a noisy and reverberant acoustic environment with $S$ static sources, where $\mathbf{p}_{s}\in\mathbb{R}^{P}$ denotes the position of the $s$-th source in the absolute coordinate system with $P$-dimensional canonical basis vectors $\mathbf{e}_{1},\dots,\mathbf{e}_{P}$ ($1\leq P\leq 3$). We assume the number of sources $S$ to be known. We consider an array with $M$ microphones whose positions are assumed to be known, with $M>P$, where $\mathbf{m}_{m}\in\mathbb{R}^{P}$ denotes the position of the $m$-th microphone. The $P\times M$-dimensional microphone positions matrix is defined as
$\mathbf{M}\;=\;[\mathbf{m}_{1},\dots,\mathbf{m}_{M}]\;.$ (1)
The origin of the coordinate system is defined at the centroid of the microphone positions, i.e., $\frac{1}{M}\mathbf{M}\mathbf{1}_{M}=\mathbf{0}_{P}$, with $\mathbf{1}_{M}$ denoting an $M$-dimensional vector of ones and $\mathbf{0}_{P}$ denoting a $P$-dimensional vector of zeros. The source and microphone positions are exemplified in Fig. 1.
@@
$\mathbf{G}_{MM}\;=\;\mathbf{M}^{\textrm{T}}\mathbf{M}\;,$ (2)
$\mathbf{G}_{s}\;=\;\mathbf{P}_{s}^{\textrm{T}}\mathbf{P}_{s}\;,$ (3)
@@
$\mathbf{M}_{\textrm{r}}\;=\;\mathbf{R}^{\textrm{T}}\mathbf{M}\;,\;$ (4)
$\mathbf{P}_{\text{r}s}\;=\;\mathbf{R}^{\textrm{T}}\mathbf{P}_{s}\;=\;[\mathbf{M}_{\textrm{r}}\,,\;\mathbf{R}^{\textrm{T}}\mathbf{p}_{s}]\;,$ (5)
@@
It is important to realize that the Gram matrices in (2) and (3) are unaffected by a basis rotation and/or reflection, i.e.,
$\mathbf{G}_{MM}\;=\;\mathbf{M}_{\textrm{r}}^{\textrm{T}}\mathbf{M}_{\textrm{r}}\;,$ (6)
$\mathbf{G}_{s}\;=\;\mathbf{P}_{\textrm{r}s}^{\textrm{T}}\mathbf{P}_{\textrm{r}s}\;.$ (7)
@@
with $\tau_{m}(\mathbf{p}_{s})$ denoting the TDOA between the $m$-th microphone and the reference microphone. The $(M+1)\times(M+1)$-dimensional EDM $\mathbf{D}_{s}$ for the $s$-th source is defined as
@@
where $\mathbf{I}_{M}$ denotes the $M\times M$-dimensional identity matrix and $\mathbf{a}_{M}$ denotes the so-called centering vector. To ensure that the centroid of the relative microphone positions $\mathbf{M}_{\textrm{r}}$ is at the origin, we define the centering vector for $\mathbf{G}_{MM}$ as $\mathbf{a}_{M}=\frac{1}{M}\mathbf{1}_{M}$. This important property will be used in the EDM-based DOA estimation in Section IV.
For the EDM-based position estimation, we extend (11) to include the distances between the $s$-th source and the microphones, i.e., for Similarly as in (11), the Gram matrix for the $s$-th source $\mathbf{G}_{s}$ can be written in terms of the EDM $\mathbf{D}_{s}$ as
@@
with $\overline{\mathbf{e}}_{m}$ the $(M+1)$-dimensional selection vector, consisting of zeros except for a one in the $m$-th entry.
@@ III-A2 TDOA Estimation and Selection
Since in practice the TDOAs of the source are obviously not available, in [25] a method was proposed to incorporate TDOA estimates into the EDM-based cost function. A commonly used method to estimate TDOAs between microphone pairs is based on the generalized cross correlation with phase transform (GCC-PHAT) function [34]. The continuous-time GCC-PHAT function between the $m$-th and the reference microphone is defined as
@@
Although the PHAT weighting in (21) has been shown to improve robustness against reverberation and noise [35, 36, 37], acoustic reflections can introduce peaks in $\xi_{m}(\tau)$ that are higher than the peak corresponding to the direct source component. Therefore, it was proposed in [25] to consider $C$ candidate TDOA estimates instead of considering only the global maximum of the GCC-PHAT function as in (22). For each microphone $m=2,\dots,M$, the set of $C$ candidate TDOA estimates is denoted as $\hat{\mathcal{T}}_{m}=\{\hat{\tau}_{m}(1),\dots,\hat{\tau}_{m}(C)\}$, corresponding to the $C$ largest local maxima of the GCC-PHAT function $\xi_{m}(\tau)$. To determine which combination of candidate TDOA estimates best fits the source and microphones geometry, we consider all $Q=C^{M-1}$ possible combinations and define the combination vectors
@@
where the index $c_{m}(q)\in\{1,\dots,C\}$ refers to one of the $C$ largest local maxima of $\xi_{m}(\tau)$. For each possible combination of candidate TDOA estimates, the vector of squared distances between the source and the reference microphone is given by
@@
The optimal combination of candidate TDOA estimates is then determined by first determining the optimal distance variable for each combination, i.e.,
@@
with corresponding Gram matrix $\hat{\mathbf{G}}_{1}={\mathbf{G}}_{1}(\hat{\alpha}(\hat{q}_{1}),\hat{q}_{1})$. It should be noted that due to possible TDOA estimation errors, $J(\hat{\alpha}(\hat{q}_{1}),\hat{q}_{1})$ is not guaranteed to be 0 unlike $J({\alpha}_{\textrm{opt}})$ from (19).
@@
and the absolute source position is then computed using (15) as

### v2

@@ Abstract
A popular method to estimate the positions or directions-of-arrival (DOAs) of multiple sound sources using an array of microphones is based on steered-response power (SRP) beamforming. For a three-dimensional scenario, SRP-based methods need to jointly optimize three continuous variables for position estimation or two continuous variables for DOA estimation. This can be computationally expensive, especially when high localization accuracy is desired. In this paper, we propose novel methods for multi-source position and DOA estimation by exploiting properties of Euclidean distance matrices (EDMs) and their respective Gram matrices. All methods require estimated time-differences of arrival (TDOAs) between the microphones. In the proposed multi-source position estimation method only a single continuous variable, representing the distance between each source and a reference microphone, needs to be optimized. For each source, the optimal continuous distance variable and set of candidate TDOA estimates are determined by minimizing a cost function that is defined using the eigenvalues of the Gram matrix. The estimated relative source positions are then mapped to estimated absolute source positions by solving an orthogonal Procrustes problem for each source. The proposed multi-source DOA estimation method entirely eliminates the need for continuous variable optimization by defining a relative coordinate system per source such that one of its coordinate axes is aligned with the respective source DOA. The optimal set of candidate TDOA estimates is determined by minimizing a cost function that is defined using the eigenvalues of a rank-reduced Gram matrix. For two sources in a noisy and reverberant environment, experimental results for different source and microphone configurations with six microphones show that the proposed EDM-based method consistently outperforms the SRP-based method in terms of position and DOA estimation accuracy. Despite the fact that the computational complexities of the EDM-based methods grow combinatorially with the number of microphones and candidate TDOA estimates, they exhibit lower run times than the SRP-based methods for the considered experimental setup with a moderate number of microphones.
@@ I Introduction
Estimating the locations, i.e., positions or directions-of-arrival (DOAs), of multiple speech sources in a noisy and reverberant environment is important for several applications, such as source tracking, speech enhancement, and speaker extraction [32, 37, 3, 45]. In this paper, we consider both compact microphone arrays with relatively small inter-microphone distances as well as spatially distributed microphones with large inter-microphone distances. For compact microphone arrays the sources will be assumed to be in the far field, such that we will focus on estimating the DOA (azimuth and elevation) of each source, whereas for spatially distributed microphones we will focus on estimating the three-dimensional position of each source.
@@
Existing model-based methods for source localization fall into two main categories [32, 37]: one-step methods such as steered-response power methods [34, 12, 8] or the subspace-based multiple signal classification (MUSIC) method [42], and two-step methods, which rely on the prior estimation of variables such as time-differences of arrival (TDOAs) between microphones [27]. In addition, several learning-based methods [49, 1, 4, 24, 50, 22, 46] for source localization have recently been proposed, showing promising results. However, most learning-based methods are trained for a specific array geometry and do not support source localization with arbitrary array geometries. In this paper, we only consider model-based localization methods, which provide the flexibility to use different array geometries without the need to retrain a neural network. A popular model-based method for single-source DOA or position estimation is based on the steered-response power with phase transform (SRP-PHAT) functional [12, 10, 33, 41, 31, 19, 23, 13]. While DOA estimation requires jointly optimizing the SRP-PHAT functional in up to two continuous variables (azimuth and elevation), position estimation requires jointly optimizing the SRP-PHAT functional in up to three continuous variables (x, y, and z coordinates). To perform accurate source localization, the SRP-PHAT functional needs to be evaluated on a discrete multi-dimensional grid with a high enough resolution in each variable resulting in a high computational complexity [19, 28]. To avoid the multi-dimensional search of SRP-based methods, in [5] we proposed a method to estimate the three-dimensional position of a single source based on Euclidean distance matrices. This method minimizes a cost function defined using the eigenvalues of a Gram matrix associated with an EDM containing the inter-microphone distances and the source-microphone distances. Using estimated TDOAs, this problem can be reformulated as an optimization in only a single distance variable, where the solution represents the distance between the source and the reference microphone. Several model-based methods have also been proposed for multi-source DOA and position estimation, e.g., based on determining multiple peaks in the SRP-PHAT functional or MUSIC spectrum [42, 8, 36, 40, 20, 17] often using clustering/association of acoustic features [36, 30, 11]. It should be noted that most of these methods focus on multi-source DOA estimation and not on multi-source position estimation.
@@
In this paper, we first extend the EDM-based single-source position estimation method in [5] to multi-source position estimation, assuming the number of sources $S$ to be known. For each combination of candidate TDOA estimates, the optimal distance variable is determined by minimizing the EDM-based cost function using a one-dimensional exhaustive search. Assuming that each source position corresponds to a unique combination of TDOAs [^1] The probability that sources have the same TDOAs in all pairs of microphones is extremely low. This would only occur for a planar (or linear) microphone array with symmetric sources with respect to this plane (or line). In this paper, we focus on three-dimensional microphone configurations., the optimal set of candidate TDOA estimates and corresponding source distances are determined, thereby also solving the association of candidate TDOA estimates to sources. The estimated relative position of each source to the microphone array can then be determined from the Gram matrix with the estimated source distance. The estimated relative positions are then mapped to estimated absolute source positions by solving an orthogonal Procrustes problem for each source. As the second contribution of this paper, we propose an EDM-based method to estimate the DOAs of multiple sources relative to a compact microphone array, which does not require continuous variable optimization. We construct a rank-reduced Gram matrix by subtracting a rank-1 matrix based on the estimated TDOAs from the Gram matrix associated with the EDM containing the microphone distances. We define a cost function using the eigenvalues of this rank-reduced Gram matrix and, similarly as for multi-source position estimation, determine the optimal set of candidate TDOA estimates by considering the $S$ smallest cost function values. The DOA of each source is then estimated from the mapping between the absolute coordinate system and the relative coordinate system, which depends on the combinations of TDOA estimates corresponding to each source.
@@
Two sets of experiments are conducted for several simulated scenarios with two sources and six microphones in a room with mild background noise and reverberation. The results of the first experiment considering spatially distributed microphones show that the proposed EDM-based method outperforms the SRP-based method in estimating the positions of the sources for all considered distances between the sources and the microphone array. Similarly, the results of the second experiment considering compact microphone arrays show that the proposed EDM-based method also consistently outperforms the SRP-based method in estimating the DOAs of the sources. In terms of computational complexity, the complexities of the EDM-based methods grow combinatorially with the number of microphones and the number of candidate TDOA estimates, whereas the complexities of SRP-based methods only grow quadratically in terms of the number of microphones. At the same time, the EDM-based methods reduce the number of continuous variables that need to be jointly optimized, resulting in lower run times.
@@ II Theoretical Background
We consider a noisy and reverberant acoustic environment with $S$ static sources, where $\mathbf{p}_{s}\in\mathbb{R}^{P}$ denotes the position of the $s$-th source in the absolute coordinate system with $P$-dimensional canonical basis vectors $\mathbf{e}_{1},\dots,\mathbf{e}_{P}$ ($1\leq P\leq 3$). We assume the number of sources $S$ to be known. We consider an array with $M$ microphones whose positions are assumed to be known, with $M>P$, where $\mathbf{m}_{m}\in\mathbb{R}^{P}$ denotes the position of the $m$-th microphone. The $P\times M$-dimensional microphone positions matrix spans $P$-dimensional space and is defined as $\mathbf{M}\;=\;[\mathbf{m}_{1},\dots,\mathbf{m}_{M}]$. The origin of the coordinate system is defined at the centroid of the microphone positions, i.e., $\frac{1}{M}\mathbf{M}\mathbf{1}_{M}=\mathbf{0}_{P}$, with $\mathbf{1}_{M}$ denoting an $M$-dimensional vector of ones and $\mathbf{0}_{P}$ denoting a $P$-dimensional vector of zeros. The source and microphone positions are exemplified in Fig. 1.
@@
$\mathbf{G}_{MM}\;=\;\mathbf{M}^{\textrm{T}}\mathbf{M}\;,$ $\mathbf{G}_{s}\;=\;\mathbf{P}_{s}^{\textrm{T}}\mathbf{P}_{s}\;,$ (1)
@@
$\mathbf{M}_{\textrm{r}}\;=\;\mathbf{R}^{\textrm{T}}\mathbf{M}\;,\;$ $\mathbf{P}_{\text{r}s}\;=\;\mathbf{R}^{\textrm{T}}\mathbf{P}_{s}\;=\;[\mathbf{M}_{\textrm{r}}\,,\;\mathbf{R}^{\textrm{T}}\mathbf{p}_{s}]\;,$ (2)
@@
It is important to realize that the Gram matrices in (1) are unaffected by a basis rotation and/or reflection, i.e.,
$\mathbf{G}_{MM}\;=\;\mathbf{M}_{\textrm{r}}^{\textrm{T}}\mathbf{M}_{\textrm{r}}\;,$ $\mathbf{G}_{s}\;=\;\mathbf{P}_{\textrm{r}s}^{\textrm{T}}\mathbf{P}_{\textrm{r}s}\;.$ (3)
@@
with $d_{1s}$ denoting the distance between the reference microphone $m=1$ and the $s$-th source, and $\tau_{m}(\mathbf{p}_{s})$ denoting the TDOA between the $m$-th microphone and the reference microphone. The $(M+1)\times(M+1)$-dimensional EDM $\mathbf{D}_{s}$ for the $s$-th source is defined as
@@
where $\mathbf{I}_{M}$ denotes the $M\times M$-dimensional identity matrix and $\mathbf{a}_{M}$ denotes the so-called centering vector. To ensure that the centroid of the relative microphone positions $\mathbf{M}_{\textrm{r}}$ is at the origin, we define the centering vector for $\mathbf{G}_{MM}$ as $\mathbf{a}_{M}=\frac{1}{M}\mathbf{1}_{M}$. This important property will be used in the EDM-based DOA estimation in Section IV. Similarly as in (7), the Gram matrix for the $s$-th source $\mathbf{G}_{s}$ can be written in terms of the EDM $\mathbf{D}_{s}$ as
@@
with $\overline{\mathbf{e}}_{M+1}$ the $(M+1)$-dimensional selection vector, consisting of zeros except for a one in the $(M+1)$-th entry.
@@ III-A2 TDOA Estimation and Selection
Since in practice the TDOAs of the source are obviously not available, in [5] a method was proposed to incorporate TDOA estimates into the EDM-based cost function. A commonly used method to estimate TDOAs between microphone pairs is based on the generalized cross correlation with phase transform (GCC-PHAT) [29]. The continuous-time GCC-PHAT function between the $m$-th and the reference microphone is defined as
@@
Although the PHAT weighting in (17) has been shown to improve robustness against reverberation and noise [9, 47, 51], acoustic reflections can introduce peaks in $\xi_{m}(\tau)$ that are higher than the peak corresponding to the direct source component, which means that the direct path peak is not guaranteed to be the global maximum. Therefore, it was proposed in [5] to consider $C$ candidate TDOA estimates, corresponding to the $C$ largest local maxima, instead of considering only the global maximum of the GCC-PHAT function as in (18), to let the EDM-based method determine the combination of candidate TDOA estimates which corresponds to the most feasible 3D source position. For each microphone $m=2,\dots,M$, the set of $C$ candidate TDOA estimates is denoted as $\hat{\mathcal{T}}_{m}=\{\hat{\tau}_{m}(1),\dots,\hat{\tau}_{m}(C)\}$, corresponding to the $C$ largest local maxima of the GCC-PHAT function $\xi_{m}(\tau)$. No association between candidate TDOA estimates and the source is applied at the TDOA estimation stage. The EDM-based method associates the best combination of candidate TDOA estimates to the source by determining which of the $Q=C^{M-1}$ possible combinations of candidate TDOA estimates corresponds to the most feasible source and microphones geometry. We define the combination vectors
@@
where the index $c_{m}(q)\in\{1,\dots,C\}$ refers to one of the $C$ largest local maxima of $\xi_{m}(\tau)$. It should be noted that the number of combinations $Q$ increases exponentially with $M$, which can be computationally demanding if $M$ is large. For each possible combination of candidate TDOA estimates, the vector of squared distances between the source and the reference microphone is given by
@@
When the candidate TDOA estimates and distance variable $\alpha$ are inconsistent with a feasible 3D geometry of the source and the microphones, the Gram matrix deviates from its ideal structure, which manifests itself as nonzero eigenvalues $\lambda_{i}(\alpha,q)$, for $i>P$ and, in some cases, even as negative eigenvalues. Using absolute values in (22) ensures that the cost function penalizes the magnitudes of these deviations, regardless of whether they appear as positive or negative eigenvalues. In other words, the cost function (22) describes the degree of geometric inconsistency of the candidate solution. The optimal combination of candidate TDOAs is then estimated by first determining the optimal distance variable for each combination, i.e.,
@@
with corresponding Gram matrix $\hat{\mathbf{G}}_{1}={\mathbf{G}}_{1}(\hat{\alpha}(\hat{q}_{1}),\hat{q}_{1})$. It should be noted that due to possible TDOA estimation errors, $J(\hat{\alpha}(\hat{q}_{1}),\hat{q}_{1})$ is not guaranteed to be 0 unlike $J({\alpha}_{\textrm{opt}})$ in (15).
@@
and the absolute source position is computed using (11) as
@@
No additional regularization beyond the eigenvalue truncation in (26) is applied for solving the orthogonal Procrustes problem.